Students taking a college introductory statistics class reported spending and average of $305 on text books that quarter with a standard deviation of $90.
One student spends $287 on textbooks. What was her standard score? What percent of the students spent less than she did on textbooks that Quarter?

Answer by stanbon(75887)

Students taking a college introductory statistics class reported spending and average of $305 on text books that quarter with a standard deviation of $90.
--------------------------------------
One student spends $287 on textbooks. What was her standard score?
z(287) = (287-305)/90 = -0.2
------------------------------------------
What percent of the students spent less than she did on textbooks that Quarter?
P(x < 287) = P(z < -0.2) = normalcdf(-100,-0.2) = 0.4207 = 42.07%
========================
Cheers,
Stan H.